Circular No. 7772

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ig IN NGC 7424 R. O. Evans, Hazelbrook, New South Wales, reports his visual dis-

covery with a 0.31-m reflector on Dec. 10.43 UT of an apparent supernova (mag ∼ 14.5) located 120′′ east and 100′′ north of the nucleus of NGC 7424. The object was confirmed visually by B. White, Linden, N.S.W., on Dec. 10.45 and via CCD by C. Bembrick, Napoleon Reefs, N.S.W., on an image taken on Dec. 10.55 (mag 15.1; offset 139′′ east and 109′′ north). Bembrick noted that nothing is visible on the Digitized Sky Survey at the position he measures for SN 2001ig: α = 22^{h}57^{m}30^{s}.69, δ = −41^{o}02′25′′.9 (equinox 2000.0). An image taken a day later by Bembrick indicates that the new object had brightened to mag 14.85.

M. M. Phillips, Observatories of the Carnegie Institute of Washington (OCIW); and N. B. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), report that SN 2001ig was observed by Suntzeff and K. Krisciunas (CTIO) with the facility CCD imager on the 1.5-m CTIO telescope, yielding the following photometry: Dec. 12.1 UT, V = 14.26, B − V = −0.05; Dec. 13.0, V = 14.24. Spectroscopic observations were obtained by R. Carlberg (University of Toronto) on Dec. 11 and by M. Gladders (OCIW) and F. Barrientos (Pontificia Universidad Católica de Chile) on Dec. 13 with the 6.5-m Baade Telescope at Las Campanas Observatory; these data were reduced by T. Matheson and S. Jha (Harvard-Smithsonian Center for Astrophysics). The first of these spectra shows only a few broad absorption features at 400, 425, 460, and 525 nm, and it defies simple classification. The Dec. 13 spectrum shows a broad P-Cyg feature with a peak at ∼ 650 nm and blue-shifted absorption at 612 nm. Also visible are broad absorption features at 408, 458, and 556 nm. The general aspect of this spectrum is similar to early spectra obtained of the type-IIb supernova 1987K (see Filippenko 1988, A.J. 96, 1941), although SN 2001ig is much bluer. If this is true, the P-Cyg feature would be Hα, and the other absorptions would correspond to Hγ, Hβ, and He I 587.6-nm at expansion velocities of 15 000–20 000 km/s. However, further spectroscopy will be needed to confirm these provisional identifications.

SUPERNOVA 2001ib IN NGC 7242 Further to IAUC 7768, R. Arbour reports that he has found another

image of SN 2001ib (at about the same brightness as on Dec. 7) taken on Dec. 5.813 UT; the image is trailed due to improper tracking.

2001 December 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green